how many molecules of h2 are produced when 332 atoms of Na reacts?

1
Expert's answer
2021-11-05T01:07:42-0400

2Na + 2H2O = 2NaOH + H2

1 mole of matter contains 6.02214076 × 1023 of chemical units.

332 atoms of Na equal to: 332 / (6.02214076 × 1023) = 55 x 10-23 moles.

According to the equation, n (H2) = 1/2 x n (Na) = 1/2 x 55 x 10-23 = 27.5 x 10-23 moles.

The amount f H2 molecules equals to: 27.5 x 10-23 x 6.02214076 × 1023 = 166